# Anomaly Prediction: A Novel Approach with Explicit Delay and Horizon

### v2

@@ Abstract
Detecting anomalies in time series data is a critical challenge across various domains. Traditional methods typically focus on identifying anomalies in immediate subsequent steps, often underestimating the significance of temporal dynamics such as delay time and horizons of anomalies, which generally require extensive post-analysis. This paper introduces a novel approach for time series anomaly prediction, incorporating temporal information directly into the prediction results. We propose a new dataset specifically designed to evaluate this approach and conduct comprehensive experiments using several state-of-the-art methods. Our results demonstrate the efficacy of our approach in providing timely and accurate anomaly predictions, setting a new benchmark for future research in this field.
@@ I Introduction
Over time, methodologies in time series anomaly detection evolved significantly [6]. Initially, the field was dominated by statistical methods and simple machine learning techniques such as ARMA [7], K-Means Clustering [8], Matrix Profile [9] etc. However, with increasing data complexity, sophisticated models like deep learning approaches such as OmniAnomaly [10], TranAD [11], Anomaly Transformer [12] have become mainstream.
While the classical approach of anomaly detection is to identify anomalies as they occur or in the immediate future [13, 14], recent research has been interested in providing alerts with sufficient lead times to enhance detection capabilities [15, 16]. Despite these advances, there are still significant difficulties in predicting the temporal dynamics of anomalies, particularly in predicting the lead time (or delay time) and precise range of predicted anomalies(or horizon), as well as the cumbersomeness of training models.
Recent critiques have also highlighted issues with current research methodologies, particularly concerning the datasets used. Many datasets do not adequately reflect real-world complexities, leading to an underestimation of deep learning methods’ effectiveness [17]. In some cases, simpler models like Autoregressive Moving Average (ARMA) outperform their complex counterparts (LSTM, Transformer), suggesting the need for more complex tasks and datasets for evaluation [18].
@@ Methods
Recognizing these challenges, this paper introduces a novel approach called ”Anomaly Prediction”. Instead of merely predicting the occurrence of an anomaly or its severity, this approach predicts the distribution density of anomaly events over future time intervals, providing clear delay time and horizon information. This enhances the practical utility of anomaly detection and demonstrates deep models’ ability to recognize complex patterns in more difficult tasks.
To empirically validate our approach, we developed a synthetic dataset tailored to assess predictive models’ performance in capturing complex temporal dynamics. We propose using Wasserstein loss to model the temporal dynamics by measuring the density distributions of anomaly events. Additionally, to thoroughly evaluate the performance of models and understand the characters of the dataset, we introduce four additional metrics (Existence, Length, Lead Time, and Dice Score) for comprehensive performance evaluation (Table I).
@@
  • Application of forecastng models and Wasserstein loss to capture the distribution density of anomaly events.
  • Development of a synthetic dataset and evaluation metrics to rigorously assess models performance.
@@
In conclusion, this study pushes the boundaries of time series anomaly detection with the introduction of ”anomaly prediction,” a novel approach that integrates delay time and horizon information. Future research will focus on refining these techniques and exploring their applicability in more complex scenarios, thereby expanding the scope and impact of anomaly detection in real-world applications.
@@ II Related Works
In this section, we provide a comprehensive review of anomaly detection models, ranging from traditional methods to contemporary deep learning methods. Next, we delve into the latest developments in time series forecasting methods. In addition, we introduce related tasks that intersect with anomaly detection, such as event prediction, human behavior detection, and root cause analysis. Finally, we present anomaly prediction as an emerging field that combines elements of anomaly detection, time series forecasting, and event prediction.
@@ II-A Anomaly Detection
Anomaly detection techniques can be roughly divided into Forecasting Methods, Reconstruction Methods, Encoding Methods, Distance Methods, Distribution Methods, and Isolation Tree Methods [19]. The key idea of detecting anomaly is identifying the samples far from the majority or different from normal samples based on a distance or a metric.
@@
Encoding Methods also encode subsequences into a low-dimensional latent space, but unlike reconstruction methods, they compute the anomaly score directly from the latent space representations without attempting to reconstruct the original subsequences. The anomaly scores are derived from the encoded representations, where anomalies are identified based on their divergence from normal patterns in the latent space. This category includes algorithms such as GrammarViz [23] and Series2Graph [24].
Distance Methods utilize specialized distance metrics to compare points or subsequences within a time series. Anomalous subsequences are expected to have larger distances from clusters of normal subsequences. Examples of algorithm include LOF [25], CBLOF [26], STAMP [27].
Distribution Methods estimate the data’s distribution or fit a distribution model to it, identifying anomalies based on their probability or distance from the calculated distributions. Anomalies are typically found in the extremes or tails of the distributions, making this approach effective for identifying infrequent patterns. Key algorithms in this category are COPOD [28], HBOS [29].
Isolation Tree Methods construct an ensemble of random trees to partition the samples of a time series. Anomalies are detected based on the number of splits required to isolate a sample, with anomalous samples having shorter paths to the root of the tree [30] [31] [32].
@@ II-B Time Series Forecasting
In time series forecasting, transformer-based models have significantly evolved to address the complexities of long-range dependencies and computational efficiency. Early models like LogTrans [33] introduced convolutional self-attention to improve the efficiency of the self-attention mechanism, while Informer [34] utilized a ProbSparse self-attention mechanism to achieve $O(L\log L)$ time complexity and memory usage, making it more scalable for long sequences. Pyraformer [35] advanced this further by capturing temporal dependencies in a multi-resolution manner, optimizing both time and space complexity. PatchTST employs PropSparse attention and patches, significantly reducing computation complexity to $O(\frac{L}{S}\log(\frac{L}{S}))$, where $S$ is the patch size, thereby enhancing the model’s efficiency and ability to process long time series.
UNet-based methods, originally designed for image segmentation, have also been adapted for time series forecasting due to their powerful feature extraction capabilities. U-Net’s architecture, with its encoder-decoder structure and skip connections, facilitates the capture and merging of multi-scale information. In the context of time series, models like U-time [36] apply convolutional layers for segmentation tasks, while YFormer [37] integrates transformer blocks at each layer of U-Net for forecasting tasks, effectively utilizing multi-resolution feature maps. Further advancing this field, Kernel U-Net [38] addresses specific challenges in time series forecasting by enhancing both expressiveness and computational efficiency. This model maintains the symmetric, hierarchical U-shaped structure of traditional U-Net but separates the process of partitioning input time series into patches from kernel manipulation, allowing for customized kernel execution, and making it more suitable for diverse and large-scale time series applications.
@@ II-C Event Prediction
In time series event prediction, methods are categorized based on their time granularity into three types: occurrence prediction, discrete-time prediction, and continuous-time prediction [39].
Occurrence prediction, the most straightforward, focuses on determining whether an event will occur within a future time period, often treated as a binary classification problem. Techniques for this include Support Vector Machines (SVMs) [40], neural networks[41], and decision trees[42]. Additionally, anomaly detection methods like one-class classification and hypothesis testing are used for rare event occurrences, leveraging the identification of deviations from normal samples. Regression-based approaches extend binary prediction to predict event count[43] or scale[44], enhancing the granularity of occurrence predictions.
Discrete-time prediction aims to forecast the approximate time slot (e.g., day, week, month) of an event’s occurrence. This approach can be direct, where future time is partitioned into discrete values and predicted using regression or ordinal classification methods[45], or indirect, where time series forecasting techniques predict future values, and events are then identified within predictions [46].
Continuous-time prediction, addressing the limitations of discrete-time methods, involves predicting the exact time of an event using techniques like regression, point processes, or survival analysis[47].
Point processes model the conditional intensity function to predict event times accurately, while survival analysis uses hazard functions to estimate event occurrence probabilities over time. These methods provide high-resolution and precise event timing, adapting to complex temporal patterns in time series data[48].
@@ II-D Early recognition of ongoing human actions
Early recognition of ongoing human actions from video streams is a critical task with broad applications across various domains, including robotics, entertainment, surveillance, and healthcare [49]. The ability to promptly detect and recognize actions as they unfold is essential for enabling real-time decision-making and enhancing operational efficiency in these fields. Recent research [49] has proposed innovative methods utilizing Bidirectional Long Short-Term Memory (BLSTM) networks to predict the temporal initiation of actions with high precision. These methods leverage advanced training techniques such as novel loss functions based on cumulative distribution functions, aiming to minimize prediction errors and improve overall recognition accuracy. By refining the estimation of action start times, these approaches demonstrate considerable enhancements in the performance of early recognition systems, showcasing their potential to advance the state-of-the-art in action detection technologies.
@@ II-E Explainable Time Series Anomaly Detection
Exathlon, a pioneering public benchmark specifically designed for explainable anomaly detection over high-dimensional time series data [50]. Exathlon features real data traces derived from extensive executions of large-scale stream processing tasks on an Apache Spark cluster. They proposed Metrics such as Anomaly Existence, Range Detection, Early Detection, and Exactly-Once Detection for unsupervised anomaly detection at instant. This benchmark enables the development and evaluation of a wide range of anomaly detection and explanation discovery techniques. In their study, the authors demonstrate the practical utility of Exathlon’s dataset, evaluation methodology, and data science pipeline design through experiments with three state-of-the-art anomaly detection and explanation methods
@@ III Method
In this section, we introduce our anomaly prediction approach, starting with problem formulation, where we define the task as predicting the distribution density of anomalies. We employ the Wasserstein distance (or cumulative sum loss) as a loss function to measure the distribution dissimilarity between actual anomalies and predicted data. To evaluate model performance, we use several metrics: the Existence of anomalies, Densities of anomalies, Lead time, and the Dice Score, which measures the overlap between predicted and actual anomalies. Our approach incorporates deep learning time series forecasting models to capture complex patterns and temporal shifts.
@@ III-A1 Anomaly Prediction Task
Consider a multivariate time series dataset represented by the matrix $x\in\mathbb{R}^{N\times M}$, where $N$ represents the number of sampling times, and $M$ represents the number of features. Let $L$ be the length of the look-back window (memory), so the slice $(x_{t+1,1},...,x_{t+L,M})$ (or, for short, $(x_{t+1},...,x_{t+L})$ contains historical information about the system at instant $t$.
Define the trajectory matrix $X_{t}\in\mathbb{R}^{L\times M}$ as a historical data segment at instant $t\in[0,N-L-1]$. The task is to use the historical data series $(x_{t+1},...,x_{t+L})$ of length $L$ to predict the probability of anomalies occurring within the next $T$ time steps.
@@
$(\hat{y}_{t+L+1},...,\hat{y}_{t+L+T})=f(x_{t+1},...,x_{t+L})$
where $f$ is a function that predicts the probabilities $\hat{Y}_{t+L+1}=(\hat{y}_{t+L+1},...,\hat{y}_{t+L+T})$ of finding anomalies based on the historical series $(x_{t+1},...,x_{t+L})$. Here, $\hat{y}_{t+L+i}$ represents the predicted probability of an anomaly at time $t+L+i$ for $i\in\{1,...,T\}$ (Figure 1).
Remark that classic time series forecasting task takes Mean Absolute Error (MAE) or Mean Square Error (MSE) as loss function. Instead of forecasting the actual value of time series, anomaly prediction predicts the distribution of anomalous events. In this case, classical MAE or MSE loss is inappropriate because the presence and absence of anomaly events with different lead time contradict each other and makes the model collapse. Therefore, We propose using Wasserstein loss, which models anomaly events as distributions over time dimension.
@@ III-A2 Wasserstein Loss
Wasserstein Loss (Cumulative Sum Loss) measures the distance or similarity of two distributions. It evaluates the overall difference between across the entire future horizon by measuring the distributional differences between predicted and actual values. Technically, this method computes the MAE or MSE of the cumulative sum loss over the time dimension. The Wasserstein loss is defined as:
$\text{Wasserstein}(\hat{Y}_{t},Y_{t})=\frac{2}{T(T+1)}\sum_{i=1}^{T}\left|\sum_{j=1}^{i}(\hat{y}_{t+j}-y_{t+j})\right|$
where the cumulative sums of the MAE of predicted values $\hat{y}$ and the true values $y$ are compared at each time step $i$ within the prediction horizon. This loss function captures both the location and the shape of the predicted distribution relative to the actual distribution. It can approximate a mixture of up to $T$ distribution of anomalies in the prediction horizon.
@@ III-B1 Existence of Anomaly
$\text{TP}(\hat{Y}_{t},Y_{t})=\begin{cases}1&\text{if }Sum(\hat{Y}_{t})>0\text{and }Sum(Y_{t})>0\\ 0&\text{otherwise}\end{cases}$
$\text{FP}(\hat{Y}_{t},Y_{t})=\begin{cases}1&\text{if }Sum(\hat{Y}_{t})>0\text{and }Sum(Y_{t})=0\\ 0&\text{otherwise}\end{cases}$
$\text{FN}(\hat{Y}_{t},Y_{t})=\begin{cases}1&\text{if }Sum(\hat{Y_{t}})=0\text{and }Sum(Y_{t})>0\\ 0&\text{otherwise}\end{cases}$
where $Sum(Y_{t})=\sum_{i=1}^{T}y_{t+i}$.
@@
$\text{Exist}(\hat{Y},Y)=\frac{2\cdot TP}{2\cdot TP+FP+FN}$
@@ III-B2 Density of Anomalies
This metric measures the difference in the cumulative densities of predicted anomalies compared to the actual count of anomaly label. It is calculated as:
$\text{Density}(\hat{Y}_{t},Y_{t})=1-\frac{1}{T}\left|\sum_{i=1}^{T}(\hat{y}_{t+i}-y_{t+i})\right|$
This metric evaluates the discrepancy in the cumulative density of predicted anomalies versus the count of actual anomaly labels. This metric ranges from 0 to 1, where 1 indicates that the size of predicted anomaly is exact, and values closer to 0 indicate over-predicting or under-predicting the size of anomaly events.

### v3

@@ Abstract
Anomaly detection in time series data is a critical challenge across various domains. Traditional methods typically focus on identifying anomalies in immediate subsequent steps, often underestimating the significance of temporal dynamics such as delay time and horizons of anomalies, which generally require extensive post-analysis. This paper introduces a novel approach for time series anomaly prediction, incorporating temporal information directly into the prediction results. We propose a new dataset specifically designed to evaluate this approach and conduct comprehensive experiments using several state-of-the-art methods. Our results demonstrate the efficacy of our approach in providing timely and accurate anomaly predictions, setting a new benchmark for future research in this field.
@@ I Introduction
Over time, methodologies in time series anomaly detection evolved significantly [6]. Initially, the field was dominated by statistical methods and simple machine learning techniques such as Autoregressive Moving Average (ARMA) [7], K-Means Clustering [8], Matrix Profile [9] etc. However, with increasing data complexity, sophisticated models like deep learning approaches such as OmniAnomaly [10], TranAD [11], Anomaly Transformer [12] have become mainstream.
While the classical approach of anomaly detection is to identify anomalies as they occur or in the immediate future [13] [14], recent research has been interested in providing alerts with sufficient lead times to enhance detection capabilities [15] [16]. Despite these advances, there are still significant difficulties in predicting the temporal dynamics of anomalies, particularly in predicting the lead time (or delay time) and precise range of predicted anomalies(or horizon), as well as the cumbersomeness of training models.
Recent critiques have also highlighted issues with current research methodologies, particularly concerning the datasets used. Many datasets do not adequately reflect real-world complexities, leading to an underestimation of deep learning methods’ effectiveness [17]. In some cases, simpler models like ARMA outperform their complex counterparts (Long Short-term Memory (LSTM), Transformer), suggesting the need for more complicated tasks and datasets [18].
@@ Methods
Recognizing these challenges, this paper introduces a novel approach called ”Anomaly Prediction”. Instead of merely predicting the occurrence of an anomaly or its severity, this approach predicts the distribution densities of anomaly events over future time intervals, providing clear delay time and horizon information. This enhances the practical utility of anomaly detection and demonstrates deep models’ ability to recognize complex patterns in more difficult tasks.
To empirically validate our approach, we develop a synthetic dataset tailored to assess predictive models’ performance in capturing complex temporal dynamics. We propose using Wasserstein loss to model the temporal dynamics by measuring the distribution densities of anomaly events. Additionally, to thoroughly evaluate the performance of models and understand the characters of the dataset, we introduce four additional metrics (Existence, Density, Lead Time, and Dice Score) for comprehensive performance evaluation (Table I).
@@
  • Application of forecasting models and Wasserstein loss to capture the distribution densities of anomaly events.
  • Development of a synthetic dataset and evaluation metrics to rigorously assess model performance.
@@
In conclusion, this study pushes the boundaries of time series anomaly detection with the introduction of ”Anomaly Prediction”, a novel approach that integrates delay time and horizon information. Future research will focus on refining these techniques and exploring their applicability in more complex scenarios, thereby expanding the scope and impact of anomaly detection in real-world applications.
@@ II Related Works
In this section, we provide a comprehensive review of anomaly detection models, ranging from traditional methods to contemporary deep learning methods. Next, we delve into the latest developments in time series forecasting methods. In addition, we introduce related tasks that intersect with anomaly detection, such as event prediction, human behavior detection, and explainable time series anomaly detection. Finally, we present anomaly prediction as an emerging field that combines elements of anomaly detection, time series forecasting, and event prediction.
@@ II-A Anomaly Detection
Supervised Anomaly Detection is similar to imbalanced binary classification. The challenges are generally from the discrimination to minority important samples and difficulty in collecting sufficient labels. [19] Unsupervised Anomaly Detection relieves these issues by taking advantage of them. The key idea of this approach is identifying the samples far from the majority or different from normal instances as anomaly. The mainstream techniques can be roughly divided into Forecasting Methods, Reconstruction Methods, Encoding Methods, Distance Methods, Distribution Methods, and Isolation Tree Methods [19].
@@
Encoding Methods also encode subsequences into a low-dimensional latent space, but unlike reconstruction methods, they compute the anomaly score directly from the latent space representations without attempting to reconstruct the original subsequences. This category includes algorithms such as GrammarViz [23] and Series2Graph [24].
Distance Methods utilize specialized distance metrics to compare points or subsequences within a time series. Anomalous subsequences are expected to have larger distances from clusters of normal subsequences [25].
Distribution Methods fit a distribution model with a given dataset and identify anomalies based on their probability or distance from the calculated distributions. Anomalies are typically found in the extremes or tails of the distributions, making this approach effective for identifying infrequent patterns. Key algorithms in this category are COPOD [26], HBOS [27].
Isolation Tree Methods construct an ensemble of random trees to partition the samples of a time series. Anomalies are detected based on the number of splits required to isolate a sample, with anomalous samples having shorter paths to the root of the tree [28].
@@ II-B Time Series Forecasting
In time series forecasting, transformer-based models have significantly evolved to address the complexities of long-range dependencies and computational efficiency. Models like LogTrans [29], Informer [30] and PatchTST reduce computation complexity from $O(L\log(L)^{2})$ to $O(\frac{L}{S}\log(\frac{L}{S}))$, where $S$ is the patch size, thereby enhancing the model’s efficiency and ability to process long time series.
UNet-based methods, originally designed for image segmentation[31], have also been adapted for time series forecasting due to their powerful feature extraction capabilities [32] [33]. U-Net’s architecture, with its encoder-decoder structure and skip connections, facilitates the capture and merging of multi-scale information. Recently, Kernel U-Net [34] has addressed specific challenges in time series forecasting by enhancing both expressiveness and computational efficiency.
@@ II-C Event Prediction
In time series event prediction, methods are categorized into occurrence prediction, discrete-time prediction, continuous-time prediction, and point process [35].
Occurrence prediction, the most straightforward, focuses on determining whether an event will occur within a future time period, often treated as a binary classification problem. Techniques for this include Support Vector Machines (SVMs) [36], and decision trees[37]. Additionally, anomaly detection methods like one-class classification and hypothesis testing are used for rare event occurrences, leveraging the identification of deviations from normal samples. Regression-based approaches extend binary prediction to predict event count[38] or scale[39], enhancing the granularity of occurrence predictions.
Discrete-time prediction aims to forecast the approximate time slot (e.g., day, week, month) of an event’s occurrence. This approach can be direct, where future time is partitioned into discrete values and predicted using regression or ordinal classification methods[40], or indirect, where time series forecasting techniques predict future feature values, and events are then identified within these predictions [41].
Continuous-time prediction, addressing the limitations of discrete-time methods, involves predicting the exact time of an event using techniques like regression, point processes, or survival analysis[42]. Point processes model the conditional intensity function to predict event time values accurately [43], while survival analysis uses hazard functions to estimate event occurrence probabilities in the next time interval [44]. These methods provide high-resolution and precise event timing, adapting to complex temporal patterns in time series data.
@@ II-D Early recognition of ongoing human actions
Early recognition of ongoing human actions from video streams is a critical task with broad applications across various domains, including robotics, entertainment, surveillance, and healthcare [45]. The ability to promptly detect and recognize actions as they unfold is essential for enabling real-time decision-making and enhancing operational efficiency in these fields. Recent research [45] has proposed innovative methods utilizing Bidirectional Long Short-term Memory (BLSTM) networks to predict the temporal initiation of actions with high precision. These methods leverage advanced training techniques such as novel loss functions based on cumulative distribution functions, aiming to minimize prediction errors and improve overall recognition accuracy.
@@ II-E Explainable Time Series Anomaly Detection
Exathlon is a benchmark specifically designed for explainable anomaly detection over high-dimensional time series data [46]. Exathlon features real data traces derived from extensive executions of large-scale stream processing tasks on an Apache Spark cluster. They propose Metrics such as Anomaly Existence, Range Detection, Early Detection, and Exactly-Once Detection for unsupervised anomaly detection at instant. This benchmark enables the development and evaluation of a wide range of anomaly detection and explanation discovery techniques.
@@ III Method
In this section, we introduce our anomaly prediction approach, starting with problem formulation, where we define the task as predicting the distribution density of anomalies. We employ the Wasserstein distance as a loss function to measure the distribution similarity between actual anomalies and predicted data. To evaluate model performance, we use several metrics related to anomalies: Existence, Densities, Lead Time, and Dice Score. Our approach incorporates deep learning time series forecasting models to capture complex patterns and temporal shifts.
@@ III-A1 Anomaly Prediction Task
Consider a multivariate time series dataset represented by the matrix $x\in\mathbb{R}^{N\times M}$, where $N$ represents the number of sampling times, and $M$ represents the number of features. Let $L$ be the length of the look-back window (memory), so the slice $(x_{t+1,1},\dots,x_{t+L,M})$ (or, for short, $(x_{t+1},\dots,x_{t+L})$ contains historical information about the system at instant $t+1$. We note $(y_{t+L+1},\dots,y_{t+L+T})$ the anomaly labels in the next $T$ step in the future horizon.
@@
$(\hat{y}_{t+L+1},\dots,\hat{y}_{t+L+T})=f(x_{t+1},\dots,x_{t+L})$ (1)
where $f$ is a function that predicts the probabilities $\hat{Y}_{t+L+1}=(\hat{y}_{t+L+1},\dots,\hat{y}_{t+L+T})$ of finding anomalies based on the historical series $(x_{t+1},\dots,x_{t+L})$. Here, $\hat{y}_{t+L+i}$ represents the predicted overlapped probabilities of anomalies at time $t+L+i$ for $i\in\{1,\dots,T\}$ (Figure 1).
Remark that classic time series forecasting task takes Mean Absolute Error (MAE) or Mean Square Error (MSE) as loss function. Instead of forecasting the actual value of time series, anomaly prediction predicts the densities of anomalous events. In this case, classical MAE or MSE loss is inappropriate because the presence and absence of anomaly events with different lead times contradict each other and make the model collapse. Therefore, We propose using Wasserstein loss, which integrates anomaly events densities over time dimension.
@@ III-A2 Wasserstein loss
Wasserstein loss (Cumulative sum loss) measures the distance or similarity of two distributions. We use it to evaluate the overall difference between predicted densities and actual Dirac distributions of anomaly labels over the entire future horizon. Technically, this method computes the MAE loss of the cumulative sum of predicted distributions and Dirac distributions over the time dimension. The Wasserstein loss is defined as:
$\text{Wasserstein}(\hat{Y}_{t},Y_{t})=\frac{2}{T(T+1)}\sum_{i=1}^{T}\left|\sum_{j=1}^{i}(\hat{y}_{t+j}-y_{t+j})\right|$ (2)
where the MAE of cumulative sums between predicted values $\hat{y}$ and true values $y$ are compared at each time step $i$ within the prediction horizon. This loss function captures both the location and the shape of the predicted distribution relative to the actual distribution. It can approximate a mixture of up to $T$ distribution of anomalies in the prediction horizon. In some cases, the coefficient $\frac{2}{T+1}$ can be ignored to simplify the calculation.
@@ III-B1 Existence of Anomaly
$\text{TP}(\hat{Y}_{t},Y_{t})=\begin{cases}1&\text{if }Sum(\hat{Y}_{t})\geq s\text{ and }Sum(Y_{t})\geq s\\ 0&\text{otherwise}\end{cases}$ (3)
$\text{FP}(\hat{Y}_{t},Y_{t})=\begin{cases}1&\text{if }Sum(\hat{Y}_{t})\geq s\text{ and }Sum(Y_{t})<s\\ 0&\text{otherwise}\end{cases}$ (4)
$\text{FN}(\hat{Y}_{t},Y_{t})=\begin{cases}1&\text{if }Sum(\hat{Y_{t}})<s\text{ and }Sum(Y_{t})\geq s\\ 0&\text{otherwise}\end{cases}$ (5)
where $Sum(Y_{t})=\sum_{i=1}^{T}y_{t+i}$ and $s\in[0,1)$ is a threshold.
@@
$\text{Exist}(\hat{Y},Y)=\frac{2\cdot TP}{2\cdot TP+FP+FN}$ (6)
@@ III-B2 Density of Anomalies
This metric measures the difference in the cumulative densities of predicted anomalies compared to the actual count of anomaly labels. It’s calculated as:
$\text{Density}(\hat{Y}_{t},Y_{t})=1-\frac{1}{T}\left|\sum_{i=1}^{T}(\hat{y}_{t+i}-y_{t+i})\right|$ (7)
This metric evaluates the discrepancy in the cumulative density of predicted anomalies versus the count of actual anomaly labels. This metric ranges from 00 to $1$, where $1$ indicates that the size of the predicted anomaly is exact, and values closer to 00 indicate over-predicting or under-predicting the size of anomaly events.